OpenStudy (anonymous):

I tried to do this question but I couldnt do it. When cooked, 85% of the kernels in a bag of microwave popcorn popped. There were 120 kernels of popcorn in the bag. How many kernels popped?

sammixboo (sammixboo):

Use the formula I showed you before =) \(\rm \dfrac{part}{whole}=\dfrac{percent}{100}\) We know what the part is 120 and the percent is 85, so let's plug in the values \(\rm \dfrac{120}{whole}=\dfrac{85}{100}\) Do you know how to solve this?

OpenStudy (anonymous):

85/100?

OpenStudy (anonymous):

100/85??

sammixboo (sammixboo):

Wait, sorry! I plugged in the info wrong xD Give me a sec

sammixboo (sammixboo):

OK so we have this formula again :D \(\rm \dfrac{part}{whole}=\dfrac{percent}{100}\) We know what the WHOLE is 120 and the percent is 85, so let's plug in the values \(\rm \dfrac{part}{120}=\dfrac{85}{100}\) NOW do you know how to solve this?

OpenStudy (anonymous):

omg hello?

sammixboo (sammixboo):

To solve this, cross multiply 120 by 85 first, so what is 120*85?

OpenStudy (anonymous):

oho kay

OpenStudy (anonymous):

10,200

sammixboo (sammixboo):

Correct, now we divide that by 100 to get what?

OpenStudy (anonymous):

102?

sammixboo (sammixboo):

Correct, you can check by dividing 102/120 and if it equals 85% (0.85) that is your answer

OpenStudy (anonymous):

102

OpenStudy (anonymous):

okay thankyou

OpenStudy (anonymous):

that was the last one thankyou sooo much